OpenStudy (anonymous):

a. find an equation for f^-1(x), the inverse function b. Verify that your equation is correct by showing that f(f^-1(x))=x and f^-1(f(x))=x f(x)=3x-1 f(x)=2/x f(x)=2x-3/x+1

OpenStudy (anonymous):

first one says multiply by 3, subtract 1. inverse will say add one divide by 3 \[f^{-1}(x)=\frac{x+1}{3}\]

OpenStudy (anonymous):

thank u soo muchh

OpenStudy (anonymous):

second one solve \[x=\frac{2}{y}\] for y get \[\frac{1}{x}=\frac{y}{2}\] or \[y=\frac{2}{x}\] or \[f^{-1}=\frac{2}{x}\] in other words it is its own inverse

OpenStudy (anonymous):

third one is the hardest, solving \[x=\frac{2y-3}{y+1}\] for that is what you need to do though